# GLUE: Gradient-free Learning to Unify Experts

### v1

@@ 4.2 Variance of two-point estimates
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=\left(\frac{{\mathcal{L}}(\boldsymbol{\alpha}+\mu\mathbf{u})-{\mathcal{L}}(\boldsymbol{\alpha}-\mu\mathbf{u})}{2\mu}\right)\mathbf{u},\;\;\mu>0$ (12)
Let $\mathbf{g}=\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ denote the true gradient of the loss with respect to $\boldsymbol{\alpha}$. As $\mu\rightarrow 0$, $\widehat{\mathbf{g}}$ becomes the directional derivative of $\mathcal{L}$ with respect to $\mathbf{u}$ and the approximation error scales with the square of $\mu$: $\widehat{\mathbf{g}}=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$. Since the directional derivative is an unbiased estimator of $g$, both $\widehat{\mathbf{g}}$ and $\overline{\mathbf{g}}$ are also unbiased [s2013stochastic]. Therefore, we bound the variance of $\overline{\mathbf{g}}$ in the following proposition—where the variance refers to the trace of the covariance matrix of the averaged gradient estimator $\overline{\mathbf{g}}$.
@@
$\mathbb{E}\left[\|\overline{\mathbf{g}}-\mathbf{g}\|^{2}\right]$ $\leq\frac{K-1}{mK}\sigma_{\mathrm{max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}}\mathcal{L}\|^{2}+O\left(\mu^{2}\right)$ (13)
@@
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$ (14)
@@
$\mathbb{E}\left[\|\widehat{\mathbf{g}}-\mathbf{g}\|^{2}\right]$ $=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})-\mathbf{g}\right\|^{2}\right]$ (15)
$=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}\right\|^{2}-2(\mathbf{g}^{\top}\mathbf{u})^{2}\right]+\|\mathbf{g}\|^{2}+O(\mu^{2})$ (16)
$=-\mathbf{g}^{\top}\mathbb{E}\left[\mathbf{u}\mathbf{u}^{\top}\right]\mathbf{g}+\|\mathbf{g}\|^{2}+O(\mu^{2})$ (17)
$=\frac{K-1}{K}\|\mathbf{g}\|^{2}+O(\mu^{2})$ (18)
@@
$\|\mathbf{g}\|^{2}=\|\nabla_{\boldsymbol{\alpha}}\mathcal{L}\|^{2}=\|\boldsymbol{\Theta}\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}\leq\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}$ (20)
@@
The proposition demonstrates that if $\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})$ is small, e.g., the expert parameters are more aligned, the variance grows slower in the number of experts $K$. The variance can be reduced with additional random direction samples (increasing $m$). Additionally assuming Lipschitz smoothness in the loss function $\mathcal{L}$ with respect to $\theta$ can be used to further upper bound $\|\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})\|^{2}$.
Fig. 1: Fine-tuning test accuracy vs. lightweight finetuning epochs (with fixed $\boldsymbol{\alpha}$). Datasets: (a) CIFAR-10, (b) SVHN, (c) Imagenette. We compare Configs 1-4 (see Sec. 5). GLUE (Config 4; ours) shows similar trends to Config 3 and consistently achieves higher test accuracy and faster convergence than Configs 1-2.
@@ 5 Experimental Evaluations
We evaluate GLUE on CIFAR-10, SVHN, and Imagenette using three backbones: ResNet-20, MobileNetV2, and an 8-layer ViT (patch size 8) respectively. We train $K=10$ expert models on heterogeneous splits of the base dataset. We adopt a non-IID experimental evaluation setup commonly used for federated learning: we sample per-class proportions for each expert from a Dirichlet distribution with concentration parameter $\delta=0.5$, then subsample a total of 2000 images to match those class proportions. In general, $\delta\leq 0.5$ induces skewed (non-IID) splits; larger $\delta$ approaches IID. Each expert is trained for 40 epochs with a batch size of 64 and an Adam optimizer with a learning rate of 0.001 and momentum parameters $(0.9,0.999)$. After training the experts, we evaluate four methods for determining $\boldsymbol{\alpha}$, the mixture coefficients used to combine the expert models’ parameters:
@@
Configs 1 and 2 do not optimize on the target domain; Configs 3 and 4 learn $\boldsymbol{\alpha}$ using target-domain training data. In Configs 3 and 4, we use Adam optimizer with learning rate 1e-2, momentum parameters (0.9, 0.99) to update $\boldsymbol{\alpha}$. Throughout training, $\boldsymbol{\alpha}$ is constrained to lie in convex hull (i.e., $\sum_{i}\alpha_{i}=1$). Additionally, for GLUE we update $\boldsymbol{\alpha}$ by randomly sampling only one random direction $u$. We use 10000 images sampled IID from the base dataset to learn the unification parameters $\boldsymbol{\alpha}$ and finetune the models as described below. All unified models are subsequently evaluated on the original testing split of each of the base datasets.
After determining $\boldsymbol{\alpha}$, we form the unified initialization $\boldsymbol{\theta}(\boldsymbol{\alpha})$ for the target model and report: (i) prior (zero-shot) performance of $\boldsymbol{\theta}(\boldsymbol{\alpha})$ without any target training, and (ii) finetuned performance after updating all network parameters on the target training set while keeping $\boldsymbol{\alpha}$ fixed. We also report the test accuracy of the individual experts before blending in Table 1 for reference. Except for the testing accuracy of experts in Table 1, we ran the experiments with three different random seeds and report the averaged test accuracy.
@@ 6 Discussion
$T_{\text{full}}-T_{\text{SPSA}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{\boldsymbol{\alpha}}-2(F+C_{\text{mix}})$ (21)
@@
As networks grow, the backward term $(\gamma-1)F$ dominates, so the gap widens in favor of SPSA. In practice, we also amortize $C_{\text{mix}}$ by caching the two blended parameter sets across the paired forward passes.
@@
Other limitations include the convex-hull restriction (the target optimum may lie outside the experts’ span), possible misalignment across experts (objectives or preprocessing), hyperparameter sensitivity of SPSA (too large radius $\mu$ of (6) biases the estimate; too small increases noise), and variance growing with $K$. These can be mitigated by light calibration, averaging a few directions, or modest schedules, but remain trade-offs.
@@ 7 Conclusion
We presented GLUE, a forward-only approach that learns a single deployable prior by convexly mixing pretrained experts and optimizing the mixture via two-point SPSA. The method avoids backpropagation and is accompanied by cost and variance analyses that explain its efficiency and stability. Empirically, GLUE delivers consistent improvements over data-size weighting and proxy-metric baselines while remaining lightweight to deploy compared to the learned full gradient mixing baseline.

### v2

@@ 4.2 Variance of two-point estimates
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=\left(\frac{{\mathcal{L}}(\boldsymbol{\alpha}+\mu\mathbf{u})-{\mathcal{L}}(\boldsymbol{\alpha}-\mu\mathbf{u})}{2\mu}\right)\mathbf{u},\;\;\mu>0$
Let $\mathbf{g}=\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ denote the true gradient of the loss with respect to $\boldsymbol{\alpha}$. As $\mu\rightarrow 0$, $\widehat{\mathbf{g}}$ becomes the directional derivative of $\mathcal{L}$ with respect to $\mathbf{u}$ and the approximation error scales with the square of $\mu$: $\widehat{\mathbf{g}}=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$. Since the directional derivative is an unbiased estimator of $g$, both $\widehat{\mathbf{g}}$ and $\overline{\mathbf{g}}$ are also unbiased [2]. Therefore, we bound the variance of $\overline{\mathbf{g}}$ in the following proposition—where the variance refers to the trace of the covariance matrix of the averaged gradient estimator $\overline{\mathbf{g}}$.
@@
$\mathbb{E}\left[\|\overline{\mathbf{g}}-\mathbf{g}\|^{2}\right]$ $\leq\frac{K-1}{mK}\sigma_{\mathrm{max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}}\mathcal{L}\|^{2}+O\left(\mu^{2}\right)$
@@
$\widehat{\mathbf{g}}(\boldsymbol{\alpha};\mathbf{u})$ $=(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})$
@@
$\mathbb{E}$ $\left[\|\widehat{\mathbf{g}}-\mathbf{g}\|^{2}\right]=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}+O(\mu^{2})-\mathbf{g}\right\|^{2}\right]$
$=\mathbb{E}\left[\left\|(\mathbf{g}^{\top}\mathbf{u})\mathbf{u}\right\|^{2}-2(\mathbf{g}^{\top}\mathbf{u})^{2}\right]+\|\mathbf{g}\|^{2}+O(\mu^{2})$
$=-\mathbf{g}^{\top}\mathbb{E}\left[\mathbf{u}\mathbf{u}^{\top}\right]\mathbf{g}+\|\mathbf{g}\|^{2}+O(\mu^{2})=\frac{K-1}{K}\|\mathbf{g}\|^{2}+O(\mu^{2})$
@@
$\|\mathbf{g}\|^{2}=\|\nabla_{\boldsymbol{\alpha}}\mathcal{L}\|^{2}=\|\boldsymbol{\Theta}\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}\leq\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})^{2}\|\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}\|^{2}$
@@
The proposition demonstrates that if $\sigma_{\mathrm{\max}}(\boldsymbol{\Theta})$ is small, e.g., the expert parameters are more aligned, the variance grows slower in the number of experts $K$. The variance can be reduced with additional random direction samples (increasing $m$). Interestingly, we find that $m=1$ already provides stable updates for small $K$ empirically (see Section 5), which is consistent with the $1/m$ scaling.
Additionally, assuming that the loss function $\mathcal{L}$ is Lipschitz smooth with respect to $\theta$ leads to an upper bound on $\|\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})\|$.
Fig. 1: Fine-tuning test accuracy vs. lightweight finetuning epochs (with fixed $\boldsymbol{\alpha}$). Datasets: (a) CIFAR-10, (b) SVHN, (c) Imagenette. We compare Configs 1-4 (see Sec. 5). GLUE (Config 4; ours) shows similar trends to Config 3 (full backpropagation for gradient computation) and consistently achieves higher test accuracy and faster convergence than Configs 1-2 (heuristics for $\boldsymbol{\alpha}$ based on data size and proxy accuracy).
@@ 5 Experimental Evaluations
We evaluate GLUE on CIFAR-10, SVHN, and Imagenette datasets using three backbones: ResNet-20, MobileNetV2, and an 8-layer ViT (patch size 8), respectively. We train $K=10$ expert models on heterogeneous splits of the base dataset. We adopt a non-IID experimental evaluation setup commonly used for federated learning: we sample per-class proportions for each expert from a Dirichlet distribution with concentration parameter $\delta=0.5$, then subsample a total of 2000 images to match those class proportions. In general, $\delta\leq 0.5$ induces skewed (non-IID) splits; larger $\delta$ approaches IID. Each expert is trained for 40 epochs with a batch size of 64 and an Adam optimizer with a learning rate of 0.001 and momentum parameters $(0.9,0.999)$. After training the experts, we evaluate four methods for determining $\boldsymbol{\alpha}$, the mixture coefficients used to combine the expert models’ parameters:
@@
Configs 1 and 2 do not optimize on the target domain; Configs 3 and 4 learn $\boldsymbol{\alpha}$ using target-domain training data. In Configs 3 and 4, we use Adam optimizer with learning rate 1e-2, momentum parameters (0.9, 0.99) to update $\boldsymbol{\alpha}$. Throughout training, $\boldsymbol{\theta}(\boldsymbol{\alpha})$ is constrained to lie in the convex hull of experts (i.e., $\sum_{i}\alpha_{i}=1$). Additionally for GLUE, we update $\boldsymbol{\alpha}$ by randomly sampling only one random direction $u$. We use 10,000 images sampled IID from the base dataset to learn the unification parameters $\boldsymbol{\alpha}$ and finetune the models as described below. All unified models are subsequently evaluated on the original testing split of each of the base datasets.
After determining $\boldsymbol{\alpha}$, we form the unified initialization $\boldsymbol{\theta}(\boldsymbol{\alpha})$ for the target model and report: (i) prior (zero-shot) performance of $\boldsymbol{\theta}(\boldsymbol{\alpha})$ without any target training, and (ii) finetuned performance after updating all network parameters on the target training set while keeping $\boldsymbol{\alpha}$ fixed. We also report the test accuracy of the individual experts before blending in Table 1 for reference. We ran all finetuning experiments with three different random seeds and report the averaged test accuracy.
@@ 6 Discussion
$T_{\text{full}}-T_{\text{SPSA}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{\boldsymbol{\alpha}}-2(F+C_{\text{mix}})$
@@
As neural networks grow, the backward term $(\gamma-1)F$ dominates, so the gap widens in favor of SPSA. In practice, we also amortize $C_{\text{mix}}$ by caching the two blended parameter sets across the paired forward passes.
@@
Other limitations include the convex-hull restriction (the target optimum may lie outside the experts’ span), possible misalignment across experts (objectives or preprocessing), hyperparameter sensitivity of SPSA (too large radius $\mu$ of (3) biases the estimate; too small increases noise), and variance growing with $K$ regardless of classical SPSA convergence-to-stationary-point guarantees. These can be mitigated by light calibration, averaging a few directions, or modest schedules, but remain trade-offs.
@@ 7 Conclusion
We presented GLUE, a forward-only approach that learns a single deployable prior by mixing pretrained experts in a convex manner and optimizing the mixture via two-point SPSA. The method avoids backpropagation and is accompanied by cost and variance analyses that explain its efficiency and stability. Empirically, GLUE delivers consistent improvements over data-size weighting and proxy-metric baselines while remaining lightweight for deployment when compared to the learned full gradient mixing baseline.